what is 12 4/7 divided by 8 7/9? please in a fraction, not a decimal.
Found 2 solutions by umail08, richwmiller:
Answer by umail08(87)

12 4/7 ÷ 8 7/9 Change each mixed number to an improper fraction
88/7 ÷ 79/9
88/7 x 9/79 … Change the division sign to multiplication and invert fraction
(88x9)/(7x79)
792/553 …. Now reduce this to a mixed number
1 239/553

Answer by richwmiller(17219)

convert to improper fractions
12 4/7=(12*7+4)/7=88/7
8 7/9=((8*9)+7)/9=79/9
(88/7) / (79/9)=
(88/7)*(9/79)=(88*9)/(7*79)